Q) A gas is filled in a container with pressure p. if the mass of the molecules is halved then thier rms speed is doubled then the resultant pressure is?
 
Q)during an experiment an ideal gas is found to ovey the equation P^2V=constant. The gas is initially at temp T and volume V. When it expands to a volume 2V the temp becomes....?
 
Q) The bulk modulus of elasticity for a monoatomic ideal gas during isothermal process (pressure of gas is P)
 
Q) A pendulum clock keeps correct time at 0 degrees. If the coeff of expansion is alpha. Then what wil be the loss in time per day if the temp rises BY t.
 
 


Grade:11

1 Answers

Sapna
6 years ago
Q.1 P =1/3 * mn/V * v²rms So when mass is halved and speed is doubled thenResultant Pressure = 1/3 * m/2* n/V *(2Vrms)²=2P